OpenStudy (anonymous):

5. Which of the following statements is not true regarding operations with exponents? (1 point) -To divide powers with the same base, subtract the exponents. -To subtract powers with the same base, divide the exponents. -To multiply powers with the same base, add the exponents. -To raise a power to a power, multiply the exponents.

OpenStudy (anonymous):

the second one, clearly \[0=2^1 - 2^1 \neq 2^{\frac{1}{1}}=2\]

OpenStudy (anonymous):

thank you